# The Basis of Foot-Sortable Sock Orderings

### v1

@@ 5. Proof of Theorem 1
Since (1), (2), and (4), all hold and $y$ and $c$ are the top two colors in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
Defined $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$. By Claim 32, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,)at_{n}\subseteq A^{\prime}bC$. Note that, with Observation 20, it is not hard to see that $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,)at_{n}$. If $y$ is not blocked by a $d^{\prime}$-sandwich in $X$, then Lemma 25 implies that a subpattern of $X$ is in $\Gamma$. Otherwise, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$.
6. Proofs of Lemmas 26, 27, and 28
See 26
@@ 6. Proofs of Lemmas 26, 27, and 28
Assume, for a contradiction, that $X$ is a minimal counterexample. Let $aba$ be the first sandwich in $X$. Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich. That is, for all $z\in X$, there exist distinct colors $x,y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$. This implies that every color must follow the first sandwich $aba$. Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$. Let $xyx$ be the first sandwich in $X$ that blocks $z$.
@@
Then $xyx=aya$ blocks $z$, so $z=b$. Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X$ contains $abayaba\cong abacaba\cong\mathcal{T}_{1}$. Otherwise, because $aba$ (and not $aya$) is the first sandwich, $X$ must contain $abya$ which is then followed by $ba$ and the color $y$. Therefore, $X$ contains either $abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $abyabay\cong abcabac\cong\mathcal{T}_{5}$.
@@
Then $xyx=byb$ blocks $z$, so $z=a$. Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$. Then $X$ contains $abaybab\cong abacbab\cong\mathcal{T}_{2}$. Otherwise, $X$ contains $abya$. Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the pattern $bab$ must follow this copy of $abya$. Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$. So, as in the previous case, $abya$ is followed by $ba$ and the color $y$. Therefore, as before, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
@@
Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$. Therefore, $X$ contains either $xabaxab\cong abcbabc\cong\mathcal{T}_{6}$ or $xabaxba\cong abcbacb\cong\mathcal{T}_{7}$.
@@
Note that $X$ must contain either $abxayx$ or $abxyax$ because $aba$ (and not $axa$ or $xyx$) is the first sandwich in $X$. Furthermore, we can assume that both $aba$ and $xyx$ in this subordering are both the first such subordering to appear in $X$. If $z=b$, then $X$ contains $abxaxba\cong abcacba\cong\mathcal{T}_{4}$. So assume $z=a$. If $y=b$, then $X$ must contain $abxabxab$ or $abxbaxab$, but then $bxb$ precedes $xbx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$. Therefore, we can assume $y\neq b$. As such, $X$ contains either $abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or $abxyax$, so we assume $X$ contains $abxyax$. Recall that every color must follow $abxya$, so every color that is not $a$ is blocked by either $axa$ or $aya$. Furthermore, $a$ is blocked by $xyx$. Therefore every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$. Since none of these three sandwiches contains $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
Case 3.2: The first $x$ appears after the first sandwich $aba$.
So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$. Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$. Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$. Hence, $X$ either contains $abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$. ∎
See 27
@@
First note that if a strict subordering of $X$ is unsortable, then $X$ is unsortable. Therefore, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable. Assume for a contradiction that no subpattern of $X$ is in $\Gamma$. Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
Note that $a$ is the only sortable color in $abtatb$ and $abtabt$. Furthermore, $b$ is the only good sortable color in $abtbI(T)at_{n}$ (when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 20).
@@
The first conjunct is given by the hypothesis of the lemma. For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP$. Therefore, because $X-d$ is sortable by assumption, $z$ is not blocked by a sandwich consisting only of colors in $P$. ∎
@@
    $S^{\prime}$ contains one of $dabt$, $adbt$, $abdt$, or for some color $e\notin P$, $S^{\prime}$ contains $abtde$; or
@@
    $R^{\prime}$ contains either $atab$ or $tP$, or
@@
    for some color $e\notin P$, $e\in S^{\prime}$ and $R^{\prime}$ contains $eP$.
@@
Case 1: A sandwich of the form $ded$ blocks $z$. In this case, we will show that (i) holds, so suppose $dabt$, $adbt$, and $abdt$ are not contained in $S^{\prime}$. Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$. To show that (i) holds, we then only need to show that $e\notin P$. Since $ded$ blocks $z$, we have $e\neq z$. We also have $e\neq y$, because, by Claim 33 a sandwich of the form $yty$ cannot block $z$ in $X=S^{\prime}dR^{\prime}$. Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$. So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$. Then
@@
By Claim 33, the color $z$ is not blocked by a $yty$ sandwich. This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$. Since $ede$ blocks $z$, we have $e\neq z$. Therefore, $e\notin\{a,b\}$.
Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$. Recall that Claim 33 implies that a $yty$ sandwich does not block $z$ in $X$. Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$. This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Furthermore, if $P=atb$ and $eP\nsubseteq R^{\prime}$, then, regardless of $e$, we must have $atab\subseteq R^{\prime}$.
First note that if $e=t$, then condition (ii) is satisfied, so assume $e\neq t$. Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$. Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$. For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$. So $z=a$ and $y=b$. Recall that we have previously shown that $ea=ez\subseteq R^{\prime}$ and that $z=a$ cannot follow the $y=b$ in $P$. Furthermore, because $eP\nsubseteq R^{\prime}$, we have $atab\subseteq R^{\prime}$. Therefore, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$. Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$. Therefore, if we consider the pattern $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable. This contradicts the fact that every strict subordering of $X$ is sortable.
@@
In this case, we have $z=b$ and $y=a$. The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 33 implies that $b$ is not blocked by a $tt_{1}t$ sandwich. Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then, regardless of $e$, we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$. So, we have the following two cases.
Case 2.2.1: $R^{\prime}$ contains $eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$. Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds. Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$. Recall that $t_{i}=e\in S^{\prime}$, so since Claim 33 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$. Therefore,
@@
a contradiction.
Case 2.2.2: $R^{\prime}$ contains $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}$. In this case, both when $n=1$ and when $n\geq 2$, we have
$X\supseteq abtdR\subseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
@@
Suppose Claim 34(i) holds. If $dabt\subseteq S^{\prime}$, then we have $at\subseteq P$ or we have $n\geq 1$ and $I(T)at_{n}\subseteq P$. This means that $X$ contains $dabtdat\cong I_{1,0}$ or $dabtdI(T)at_{n}\cong I_{1,n}$. Otherwise, $X$ contains one of the patterns in the following table.
@@
If Claim 34(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ contains either $I_{1,n}$ or one of the patterns from the table above.
Finally, if Claim 34(ii) holds, then $X$ must contain one of the four patterns in the following table.
@@
As in the proof of Lemma 27, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$. For a contradiction, assume that no subpattern of $X$ is in $\Gamma$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$. Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$. Therefore, there exists a color that is neither $a$ or $t$ in $R$. Let $r$ be the first color that is neither $a$ nor $t$ in $R$. Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
@@
Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$. Note that then $x$ is one of the first two colors to appear in $tR$. Lemma 24 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $ax\nsubseteq R$. Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$. Since $y$ is unsortable in $(at,R)$, Lemma 24 implies that $ay\subseteq R$. Then $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$. However, with Lemma 22, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, implies that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction. ∎
@@
First, suppose that $rat\subseteq R$. Since, by Claim 35, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$. Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the pattern $ratr$ does not appear $R$, so we are done in this case.
Now, suppose $rat\nsubseteq R$. Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the sandwich that blocks $t$ in $atR$ cannot contain $a$. So there exist distinct colors $c,b$ that are not $a$ or $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$. Since, by Claim 35, we have $at\subseteq R$, and $r$ is the first color that is not $a$ or $t$ to appear in $R$, and $rat\nsubseteq R$, we have either $r=c$ and $arbrt\subseteq R$ or $r\notin\{b,c\}$ and $arcbct\subseteq R$. ∎
@@
For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$. Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$. In particular, this implies that $rt\nsubseteq S$ and $atPa\nsubseteq X$.
@@
Note that Claims 36 and 37 establish the first statement of the conclusion.
@@
    $S^{\prime}$ contains one of $dat$, $adt$, or, for some color $e\notin P$, $S^{\prime}$ contains $atde$; or
@@
    $R^{\prime}$ contains one of $taxat$, $atxat$, $txaxt$, or $taxyxt$ where $a$, $t$, $x$, and $y$ are distinct colors; or
@@
    for some color $e\notin P$, $R^{\prime}$ contains $eP$ and $S^{\prime}$ contains one of $eat$, $aet$, or $ate$

### v2

@@ 2.5. Proof of Theorem 3
Since (3), (4), and (6), all hold and $y$ and $c$ are the top two colors on $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
Define $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$. By Claim 15, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$. Note that $y$ and $c$ are the only sortable colors in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$, because every other color is blocked by a $ycy$ sandwich. Furthermore, after sorting $c$ we have $(ayyt_{0},I(t_{0},\dotsc,t_{n})at_{n})$ which Observation 2 implies is unsortable. Therefore, with Observation 2, $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$. This means that the hypothesis of Lemma 8 is satisfied with $y$, $d^{\prime}$, $ayc$, and $yI(c,t_{0},\dotsc,t_{n})at_{n}$ playing the roles of $a$, $d$, $S^{\prime}$ and $R^{\prime}$, respectively. Therefore, we can assume that $y$ is blocked by a $d^{\prime}$-sandwich in $X$. Then Lemma 10 implies that a subpattern of $X$ is in $\Gamma$.
2.6. Proofs of Lemmas 9, 10, and 11
See 9
@@ 2.6. Proofs of Lemmas 9, 10, and 11
Assume, for a contradiction, that $X$ is a minimal counterexample. Let $aba$ be the first sandwich in $X$. Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich. That is, for all $z\in X$, there exist distinct colors $x$ and $y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$. This implies that every color must follow the first sandwich $aba$. Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$. Let $xyx$ be the first sandwich in $X$ that blocks $z$.
@@
Then $xyx=aya$ blocks $z$, so $z=b$. Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X\supseteq abayaba\cong abacaba\cong\mathcal{T}_{1}$. Otherwise, because $aba$ (and not $aya$) is the first sandwich, we have $X\supseteq abya$. This $abya$ subordering is followed by $ba$ and the color $y$ in $X$. Therefore, either $X\supseteq abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $X\supseteq abyabay\cong abcabac\cong\mathcal{T}_{5}$.
@@
Then $xyx=byb$ blocks $z$, so $z=a$. Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$. Then $X\supseteq abaybab\cong abacbab\cong\mathcal{T}_{2}$. Otherwise, because $aba$ is the first sandwich (and not $aya$), we have $X\supseteq abya$. Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the subordering $bab$ must follow the subordering $abya$. Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$. So, as in the previous case, $abya$ is followed by $ba$ and the color $y$. Therefore, as in the previous case, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
@@
Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$. Therefore, $X\subseteq xabaxab\cong abcbacb\cong\mathcal{T}_{6}$ or $X\subseteq xabaxba\cong abcbabc\cong\mathcal{T}_{7}$.
@@
By the case, an $axa$ sandwich precedes the first $xyx$ sandwich. Since the first sandwich that blocks $z$ is $xyx$, we must have $z=a$. Note that, because $xyx$ blocks $z=a$, we also have $y\neq a$.
The color $x$ must follow the $ab$ in the first sandwich $aba$, as otherwise an $axa$ sandwich would proceed the first $aba$ sandwich. Since $xyx$ is not the first sandwich in $X$ and $y\neq a$, either $X\supseteq abxayx$ or $X\supseteq abxyax$. We can assume that both $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=b$, then the subordering is $abxabx$ or $abxbax$. But then $bxb$ precedes $xbx=xyx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$. Therefore, we can assume $y\neq b$. Therefore, the colors $a$ and $b$ are distinct from the colors $x$ and $y$, so either $X\supseteq abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or the subordering is $abxyax$. So, we can assume the subordering is $abxyax$.
Recall that every color must follow the first $aba$ sandwich, so every color must follow $abxya$. This with the fact that $a$ is neither $x$ nor $y$ implies that every color that is not $a$ is blocked by either $axa$ or $aya$. Furthermore, $a$ is blocked by $xyx$. Therefore, every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$. Since none of these three sandwiches includes the color $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
Case 3.3: The first $x$ appears after the first sandwich $aba$.
So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$. Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$. Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$. Hence, $X\supseteq abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $X\supseteq abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$. ∎
See 10
@@
First note that if a proper subordering of $X$ is unsortable, then $X$ is unsortable. Therefore, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable. Assume for a contradiction that no subpattern of $X$ is in $\Gamma$. Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
Note that $a$ is the only sortable color in $abtatb$ and $abtabt$. We claim that $b$ is the only good sortable color in $abtbI(T)at_{n}$. To see this note that when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 2.
@@
The first conjunct is given by the hypothesis of the lemma. For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP\subseteq X-d$. Therefore, because $X-d$ is a proper subordering of $X$ and is therefore sortable, $z$ is not blocked by a sandwich consisting only of colors in $P$. ∎
@@
    one of $dabt$, $adbt$, $abdt$, is a subordering of $S^{\prime}$ or $abtde\subseteq S^{\prime}$ for some color $e\notin P$; or
@@
    $atab\subseteq R^{\prime}$ or $tP\subseteq R^{\prime}$; or
@@
    for some color $e\notin P$, $e\in S^{\prime}$ and $eP\subseteq R^{\prime}$.
@@
Case 1: A sandwich of the form $ded$ blocks $z$. In this case, we will show that (i) holds, so suppose that neither $dabt$, $adbt$, nor $abdt$ is a subordering of $S^{\prime}$. Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$. To show that (i) holds, we then only need to show that $e\notin P$. Since $ded$ blocks $z$, we have $e\neq z$. We also have $e\neq y$, since otherwise we would have a $yty$ sandwich in $S^{\prime}$ which would block $z$ and this is forbidden by Claim 16. Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$. So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$. Then
@@
By Claim 16, the color $z$ is not blocked by a $yty$ sandwich. This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$. Since $ede$ blocks $z$, we have $e\neq z$. Therefore, $e\notin\{y,z\}=\{a,b\}$.
Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$. Recall that Claim 16 implies that a $yty$ sandwich does not block $z$ in $X$. Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$. This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P$ begins with $zy$. So, we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. If $P=atb$ and $eP\nsubseteq R^{\prime}$, then, because $ea=ez\subseteq R^{\prime}$ and $z=a$ cannot follow $y=b$ in $P$, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Note that, in both cases, we have $atab\subseteq R^{\prime}$.
First note that if $e=t$, then the preceding argument implies that condition (ii) is satisfied, so assume $e\neq t$. Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$. Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$. For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$. Recall that we have previously shown that $e\in S^{\prime}$ and that, in this case, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$. Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$. Therefore, if we consider the subordering $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable. This contradicts the fact that every proper subordering of $X$ is sortable.
@@
In this case, we have $z=b$ and $y=a$. The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 16 implies that $b$ is not blocked by a $tt_{1}t$ sandwich. Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$. But then,
$X\supseteq abtdR\supseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
a contradiction.
Therefore, we can assume that $R^{\prime}\supseteq eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$. Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds. Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$. Recall that $t_{i}=e\in S^{\prime}$, so since Claim 16 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$. Therefore,
@@
Suppose Claim 17(i) holds. Note that $at\subseteq P$ or $n\geq 1$ and $I(T)at_{n}\subseteq P$. Therefore, if $dabt\subseteq S^{\prime}$, then either $X\supseteq dabtdat\cong I_{1,0}$ or $X\supseteq dabtdI(T)at_{n}\cong I_{1,n}$. Otherwise, one of the orderings in the following table is a subordering of $X$.
@@
If Claim 17(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ either $I_{1,n}\preceq X$ or one of the orderings in the table above is a subpattern of $X$.
Finally, if Claim 17(ii) holds, then one of the four orderings in the following table is a subordering of $X$.
@@
As in the proof of Lemma 10, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$. For a contradiction, assume that no subpattern of $X$ is in $\Gamma$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$. Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$. Therefore, there exists a color that is neither $a$ nor $t$ in $R$. Let $r$ be the first color that is neither $a$ nor $t$ in $R$. Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
@@
Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$. Note that then $x$ is one of the first two colors to appear in $tR$. Lemma 7 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $x\neq a$ and $ax\nsubseteq R$. Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$. Since $y$ is unsortable in $(at,R)$, Lemma 7 implies that either $y=a$ or $ay\subseteq R$. In either case, $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$. However, with Lemma 5, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, imply that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction. ∎
@@
First, suppose that $rat\subseteq R$. Since, by Claim 18, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$. Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the subordering $ratr$ does not appear $R$, so we are done in this case.
Now, suppose $rat\nsubseteq R$. Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the first sandwich that blocks $t$ in $atR$ does not include the color $a$. So there exist distinct colors $c$ and $b$ that are both neither $a$ nor $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$. Since $r$ is the first color that is not $a$ or $t$ to appear in $R$, we have either $r=c$ and $rbrt\subseteq R$ or $r\notin\{b,c\}$ and $rcbct\subseteq R$. By Claim 18, we have $at\subseteq R$. This, with the fact that $rat\nsubseteq R$, implies that either $arbrt\subseteq R$ or $arcbct\subseteq R$, where the distinct colors $b$ and $c$ are distinct from $a$, $r$, and $t$. ∎
@@
For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$ and $a$ is a good sortable color in $atP$. Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$.
@@
Note that Claims 19 and 20 establish the first statement of the conclusion.
@@
    $dat$ or $adt$ is a subordering of $S^{\prime}$ or $atde\subseteq S^{\prime}$ for some color $e\notin P$; or
@@
    $taxat$, $atxat$, $txaxt$, or $taxyxt$ is a subordering of $R^{\prime}$ where $a$, $t$, $x$, and $y$ are distinct colors; or
@@
    $eP\subseteq R^{\prime}$ for some color $e\notin P$ and $eat$, $aet$, or $ate$ is a subordering of $S^{\prime}$.